# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

### v1

@@ 4.3 First demonstration
The aim of this first demonstration is to investigate the performance of the proposed metrics on a simple problem. At this stage, no surrogate model is used; instead, the methods are applied directly to the actual function. We conducted a comparison with first-order, second-order, and total order Sobol' indices, and averaged SHAP (referred to as $S_{F}$, $S_{T}$, and $\bar{Sh}$, respectively). For the analysis of two-way interactions, we compared the corresponding metrics for two-way interactions, denoted as $I_{pdp,ij}$, $S_{ij}$, and $\bar{Sh}_{ij}$, where $i$ and $j$ represent the indices of the variables, with $i\neq j$. Normalized Shapley effects lie between first-order and total Sobol’ indices [41]. Accordingly, Sobol’ and Shapley derived indices together provide transparent, extensible variance-based diagnostics of input effects and interactions for practical design in engineering.
@@
where $x_{i}\in[-1,1]$ for $i=1,2,3$. The main challenge with this problem is that $x_{2}$ strongly interacts with $x_{3}$ in such a way that the PDP yields a flat curve (therefore, the $I_{pdp}$ is zero or close to zero). We did not perform a thorough analysis of this problem; instead, we use this problem to check the adequacy of the new metrics (i.e., $\mu_{I,ice}$, $\sigma_{I,ice}$, and $\sigma_{\rho}$). The Sobol' indices and averaged SHAP were calculated using $10^{6}$ samples. Due to the low dimensionality of the problem, the SHAP values are calculated analytically. To calculate the PDP and ICE-based metrics, 10,000 ICE curves were drawn from the actual function but the SMT-Explainability toolbox allows us to do the same based on a surrogate model [46, 47].
@@
The centered PDP and ICE plots for this problem are shown in Fig. 1. It is evident that the influence of $x_{2}$ on $y$ varies depending on the value of $x_{3}$, and vice versa. As a result, the PDPs for $x_{2}$ and $x_{3}$ appear as almost flat lines, leading to a near-zero PDP-based feature importance for both variables. However, this is misleading, as the true importance of these variables is not zero. On the other hand, $x_{1}$ does not interact with $x_{2}$ and $x_{3}$, which is why its PDP and ICE curves are essentially identical. Moreover, its overall impact is actually less significant than that of $x_{2}$ and $x_{3}$.
@@
The complete GSA results are presented in Table 1. The first-order Sobol' indices are essentially zero for all variables. However, the total order Sobol' indices are close to 1 for $x_{2}$ and $x_{3}$, suggesting that the only contribution to the total variance comes from the interaction between $x_{2}$ and $x_{3}$. As anticipated, $I_{pdp}$ does not suggest that $x_{2}$ and $x_{3}$ are more significant than $x_{1}$, due to the previously mentioned challenge. In contrast, Averaged SHAP accurately identifies the importance of $x_{2}$ and $x_{3}$. Unlike $I_{pdp}$, $\bar{Sh}$ is based on the impact of each feature on the model’s predictions and is capable of detecting subtle yet crucial relationships that $I_{pdp}$ might overlook. It can also be observed that ICE-based feature importance yields finite and reasonable values for $x_{2}$ and $x_{3}$. Notably, $\mu_{I,ice}$ effectively identifies the importance of $x_{2}$ and $x_{3}$ by focusing on the importance of each ICE curve. Thus, it can be concluded that $\mu_{I,ice}$ addresses the main limitation of $I_{pdp}$. The $\sigma_{I,ice}$ value that equals 0 for $x_{1}$ is the consequence of its non-interacting nature. However, $\sigma_{I,ice}=0$ could also mean another thing. In fact, a zero value of $\sigma_{I_{\mathrm{ice}}}$ means the per-instance ICE spread is identical for all conditioning values; this can indicate either (i) a genuinely non-interacting additive input, as in our example, or (ii) the presence of interactions that alter curve shape (e.g. flip sign or slope) while preserving the integrated spread. Hence, in this second case, visual inspection of ICE curves or slope-based diagnostics is required to disambiguate these cases, especially when relying on surrogate models. For the case of $x_{2}$, this means that the importance magnitude of each ICE curve of $x_{2}$ is the same, although the gradient shifts depending on how it interacts with $x_{3}$. Lastly, the finite value of $\sigma_{I,ice}$ for $x_{3}$ is reasonable because, as can be observed in Fig. 1(c), each ICE curve yields different $I_{ice}$.
Lastly, $\sigma_{\rho}$ is not a measure of GSA but a change in the trend of ICE curves with respect to the PDP. The fact that $\sigma_{\rho}=0$ for $x_{1}$ suggests that $x_{1}$ does not interact with other variables. Furthermore, the values of $\sigma_{\rho}=1$ for $x_{2}$ and $x_{3}$ indicates strong impact of interactions on the trend. As suggested by Fig. 1 again, the correlations between ICE curves and PDP for $x_{2}$ are either -1 or 1. Although the finite correlations result from the calculation process and should theoretically be zero when correlating with a flat function, this nonetheless highlights a subtle shift in the trend. As for $x_{3}$, the sharp discontinuity is the reason why the standard deviation of the ICE correlation values is non-zero.
@@
$\mu_{I,ice}$ 0.116 2.908 2.505
$\sigma_{I,ice}$ 0.000 0.000 1.449
$\sigma_{\rho}$ 0.000 1.000 1.000
The two-way GSA results are presented in Table 2. It is noteworthy that all metrics (namely, two-way $I_{pdp}$, two-way $\bar{Sh}_{ij}$, and second-order Sobol' indices) successfully identify the interaction between $x_{2}$ and $x_{3}$. It is an expected behavior as Second-order Sobol' indices quantify how much the combined interaction of two input variables contributes to the total output variance, over and above the sum of their individual effects. In the case of $\mu_{I,ice}$, this interaction information is incorporated into $\mu_{I,ice}$ and complements it. Specifically, $\mu_{I,ice}$ and its corresponding $\sigma_{I,ice}$ reflect the overall impact of the interaction, while $I_{pdp,ij}$ quantifies the influence of specific interactions.
@@ 5 Demonstration on analytical and engineering problems
Metrics for analytical problems were evaluated on the original function, while surrogate models were used for engineering cases where the true function was unknown. For all problems, we conducted a comparison with other sensitivity metrics (e.g., Sobol' indices and averaged SHAP).
@@ 5.1.1 Global sensitivity analysis
First, let us compute the GSA values from Sobol' indices, averaged SHAP, PDP, and ICE. The comparison of various GSA metrics is presented in Table 3. Notably, all metrics produce the same ranking of importance, with the most significant variable being $x_{4}$ followed by $x_{1}$ and $x_{2}$, with $x_{3}$ and $x_{5}$ are considered to be less significant. Interestingly, $I_{pdp}$ and $\mu_{I,ice}$ provide identical importance values for $x_{3}$, $x_{4}$, and $x_{5}$. This occurs because the centered ICE curves for variables without interactions are essentially the PDP itself, leading to the mean of $I_{ice}$ (i.e., $I_{pdp}$) being identical to $I_{pdp}$. Conversely, the $\mu_{I,ice}$ values for $x_{1}$ and $x_{2}$ are higher than their $I_{pdp}$ counterparts, which is due to the interaction effects considered in the calculation of $\mu_{I,ice}$. In this regard, if the value $I_{pdp}$ equals $\mu_{I,ice}$, this signals that the interaction is not present. Conversely, when $\mu_{I,ice}>I_{pdp}$, this signals that the interaction term is present since this means that the ICE curves are not exactly similar to the PDP in terms of the trend and impact. Note that this approach is similar to interpreting first-order and total order Sobol' indices. It is also evident in this problem that the values of $S_{T}$ and $S_{F}$ are identical for $x_{3}$, $x_{4}$, and $x_{5}$, while $S_{T}$ for $x_{1}$ and $x_{2}$ are higher than $S_{F}$ for $x_{1}$ and $x_{2}$.
@@
The barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$ for the Friedman regression problem is further shown in Fig. 2(a). Here, it can be seen that $\mu_{I,ice}$, with the errorbar showing $\mu_{I,ice}\pm\sigma_{I,ice}$ provides more information than $I_{pdp}$ and $\bar{Sh}$. Again, a large vertical bar indicates a stronger effect of interactions, as evidenced in $x_{1}$ and $x_{2}$, but non-existent in other variables. Further, Fig. 2(b) depicts the boxplots of the correlation between the ICE curves and PDP. In this figure (and similar figures throughout the paper), the lower and the upper bounds of the box show the first ($Q_{1}$) and the third quartile ($Q_{3}$), respectively, the solid line within the box shows the median ($Q_{2}$), the lower and the upper“whisker” show $Q_{1}-1.5(Q_{3}-Q_{1})$ and $Q_{3}+1.5(Q_{3}-Q_{1})$, and the crosses show individual data that lie outside the “whiskers”. For $x_{1}$ and $x_{2}$, the overall correlations are positive, but some ICE curves deviate significantly from the PDP (indicating strong interaction), particularly those with lower correlation values. The non-zero values of $\sigma_{\rho}$ also signify that the effect of the interaction changes the trend of the ICE curves. It is worth noting that a non-zero value of the correlation between the ICE curve and PDP indicates a change in nonlinearity or trend. The $\sigma_{\rho}$ value for $x_{3}$, $x_{4}$, and $x_{5}$ is exactly zero because the trend for the ICE curve is exactly similar to its PDP. Let us now analyze the interaction GSA-values as shown in Table 4. The two-way PDP feature importance, SHAP, and second-order Sobol' indices (see Table 4). further confirm the presence of the interaction between $x_{1}$ and $x_{2}$. In this regard, Second-order Sobol indices quantify the interaction effects between pairs of input variables on the output variance, while interaction SHAP values measure the specific contribution of interactions between inputs to a model’s prediction. It was observed that the estimated interaction metrics are zero or nearly zero for all cross terms but $x_{1}-x_{2}$. Note that this information should be digested together with an analysis of the visualization of PDP and ICE, as shown next.
(a) Barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$.
(b) Boxplot of ICE-based correlation values.
Figure 2: Barplot of some GSA metrics and boxplot of ICE correlation values for the Friedman regression problem.
@@ 5.1.2 Visualization
The first variable of interest is $x_{1}$, with the centered PDP/ICE and SHAP plots shown in Figs. 3 and 4, respectively. The plots for $x_{2}$ are similar, as both variables appear only in the term $10\sin(\pi x_{1}x_{2})$. The curves are colored by $x_{2}$ and $x_{4}$, where $x_{1}$ interacts solely with $x_{2}$. Notably, the centered PDP/ICE and SHAP plots reveal distinct insights: the effect of $x_{1}$ varies with the value of $x_{2}$, reflecting their interaction. For instance, when $x_{2}$ is small, the impact of $x_{1}$ is minimal, resulting in an almost flat ICE curve. However, as $x_{2}$ increases, the impact of $x_{1}$ becomes more significant, with the overall trend first increasing and then decreasing (which is due to the sinusoidal term). This effect is quantified and shown by the non-zero values of $\sigma_{I,ice}$ and $\sigma_{\rho}$ discussed earlier. Further, it can be seen that the coloring seems to be disordered when the ICE curves for $x_{1}$ are color-coded according to $x_{4}$; which indicates that there is no interaction between how the change in $x_{4}$ affects $x_{1}$.
@@
We now address the uncertainty analysis of a wind turbine fatigue problem, where the uncertainties involve five independent random variables related to wind conditions [49]. The input variables considered are wind speed ($V_{hub}$), wind direction ($\theta_{w}$), wave height ($H_{s}$), wave period ($T_{p}$), and wind-platform direction ($\theta_{p}$) with their histograms are shown in Fig. 6. Notice that the data was normalized within $1,1]^{5}$, as provided by the original paper. The side-side tower base bending moment ($M_{x,twr}$) is selected as the output of interest due to its nonlinear behavior, evaluated using FAST, NREL’s aeroelastic code for simulating the coupled dynamic response of wind turbines [50]. Subsequently, a PCE model with $p_{max}=5$, $q=1$, and data-driven orthogonal polynomials was built using 2000 samples, yielding $R^{2}=0.97$ which is more than sufficient for knowledge discovery purposes.

### v3

@@ 4.4 First demonstration
The aim of this first demonstration is to demonstrate the proposed evaluation workflow, highlighting how visualization and ICE-based metrics are used together on a simple problem. At this stage, no surrogate model is used; instead, the methods are applied directly to the actual function. We conducted a comparison with first-order, second-order, and total order Sobol’ indices, and averaged SHAP (referred to as $S_{F}$, $S_{T}$, and $\bar{Sh}$, respectively). For the analysis of two-way interactions, we compared the corresponding metrics for two-way interactions, denoted as $I_{pdp,ij}$, $S_{ij}$, and $\bar{Sh}_{ij}$, where $i$ and $j$ represent the indices of the variables, with $i\neq j$.
@@
where $x_{i}\in[-1,1]$ for $i=1,2,3$. The main challenge with this problem is that $x_{2}$ strongly interacts with $x_{3}$ in such a way that the PDP yields a flat curve (therefore, the $I_{pdp}$ is zero or close to zero). We did not perform a thorough analysis of this problem; instead, we use this problem to check the adequacy of the new metrics (i.e., $\mu_{I_{ice}}$, $\sigma_{I_{ice}}$, and $\sigma_{\rho}$). The Sobol’ indices and averaged SHAP were calculated using $10^{6}$ samples. Due to the low dimensionality of the problem, the SHAP values are calculated analytically. To calculate the PDP and ICE-based metrics, 10,000 ICE curves were drawn from the actual function but the SMT-Explainability toolbox allows us to do the same based on a surrogate model [45, 42].
The anchored PDP and ICE plots for this problem are shown in Fig. 1. In this figure, as well as in similar figures throughout the paper, the one-dimensional PDP is represented by a dashed line, whereas the corresponding ICE curves are shown as solid lines. It is evident that the influence of $x_{2}$ on $y$ varies depending on the value of $x_{3}$, and vice versa. As a result, the PDPs for $x_{2}$ and $x_{3}$ appear as almost flat lines, leading to a near-zero PDP-based input importance for both variables. However, this is misleading, as the true importance of these variables is not zero. On the other hand, $x_{1}$ does not interact with $x_{2}$ and $x_{3}$, which is why its PDP and ICE curves are essentially identical. Moreover, its overall impact is actually less significant than that of $x_{2}$ and $x_{3}$.
@@
The complete GSA results are presented in Table 1. The first-order Sobol’ indices are very small for all variables (they should essentially be zero for $x_{2}$ and $x_{3}$ due to their flat PDPs). However, the total order Sobol’ indices are close to 1 for $x_{2}$ and $x_{3}$, suggesting that their only contribution to the total variance comes from the interaction between $x_{2}$ and $x_{3}$. As anticipated, $I_{pdp}$ does not suggest that $x_{2}$ and $x_{3}$ are more significant than $x_{1}$, due to the previously mentioned challenge. In contrast, Averaged SHAP accurately identifies the importance of $x_{2}$ and $x_{3}$. Unlike $I_{pdp}$, $\bar{Sh}$ is based on the impact of each input on the model’s predictions and is capable of detecting subtle yet crucial relationships that $I_{pdp}$ might overlook. It can also be observed that ICE-based feature importance yields finite and reasonable values for $x_{2}$ and $x_{3}$. Notably, $\mu_{I_{\mathrm{ice}}}$ effectively identifies the importance of $x_{2}$ and $x_{3}$ by focusing on the importance of each ICE curve. Thus, it can be concluded that $\mu_{I_{\mathrm{ice}}}$ addresses the main limitation of $I_{pdp}$. The $\sigma_{I_{\mathrm{ice}}}$ value that equals 0 for $x_{1}$ is the consequence of its non-interacting nature (note that we analyze the standard deviation instead of the variance of $I_{\mathrm{ice}}$). However, $\sigma_{I_{\mathrm{ice}}}=0$ can indicate either (i) a genuinely non-interacting additive input, as in the case of $x_{1}$, or (ii) the presence of interactions that alter curve shape (e.g., flip sign or slope) while preserving the integrated spread. Hence, in this second case, visual inspection of ICE curves or slope-based diagnostics is required to disambiguate these cases, especially when relying on surrogate models. Indeed, for the case of $x_{2}$, this means that the importance magnitude of each ICE curve of $x_{2}$ is the same, although the gradient shifts depending on how it interacts with $x_{3}$. Distinguishing between these two scenarios requires an explicit examination of interaction metrics, which are not captured by $\sigma_{I_{\mathrm{ice}}}$. In fact, as shown in Table 2, $x_{2}$ is found to strongly interact with $x_{3}$. Conversely, the table also shows that $x_{1}$ does not interact with other variables. Lastly, the finite value of $\sigma_{I_{\mathrm{ice}}}$ for $x_{3}$ is reasonable because, as can be observed in Fig. 1(c), each ICE curve yields a different $I_{\mathrm{ice}}$.
It is worth discussing the different rankings produced by the total Sobol’ index $S_{T}$ and the proposed ICE-based importance metric $\mu_{I_{\mathrm{ice}}}$, as illustrated in this problem. The total Sobol’ is a GSA index that assigns nearly identical importance to $x_{2}$ and $x_{3}$, reflecting the fact that the dominant contribution to the output variance arises from their interaction, which is fully attributed to both variables in the total-effect formulation. In contrast, $\mu_{I_{\mathrm{ice}}}$ is an averaged local index that ranks $x_{2}$ as more important than $x_{3}$, since it measures the mean conditional variability along ICE curves and therefore captures how strongly each variable influences the model response across different conditional settings. This example highlights that while $S_{T}$ quantifies variance contribution, $\mu_{I_{\mathrm{ice}}}$ characterizes conditional response behavior, leading to different but complementary importance rankings.
Lastly, $\sigma_{\rho}$ is not a measure of GSA but a change in the trend of ICE curves with respect to the PDP. The fact that $\sigma_{\rho}=0$ for $x_{1}$ suggests that $x_{1}$ does not interact with other variables. Furthermore, the values of $\sigma_{\rho}$ are not calculated for $x_{2}$ and $x_{3}$ because it does not satisfy the criterion mentioned in Eq. (22). In particular, this signals a flat PDP and a very strong interaction. Indeed, as can be seen in Fig. 1, the PDPs for $x_{2}$ and $x_{3}$ are essentially flat, while the ICE curves indicate a significant change of trends.
@@
$\mu_{I_{\mathrm{ice}}}$ 0.116 2.908 2.505
$\sigma_{I_{\mathrm{ice}}}$ 0.000 0.000 1.449
$\sigma_{\rho}$ 0.000 - -
The two-way GSA results are presented in Table 2. It is noteworthy that all metrics (namely, two-way $I_{pdp}$, two-way $\bar{Sh}_{ij}$, and second-order Sobol’ indices) successfully identify the interaction between $x_{2}$ and $x_{3}$. It is an expected behavior as Second-order Sobol’ indices quantify how much the combined interaction of two input variables contributes to the total output variance, over and above the sum of their individual effects. In the case of $\mu_{I_{\mathrm{ice}}}$, this interaction information is incorporated into $\mu_{I_{\mathrm{ice}}}$ and complements it. Specifically, $\mu_{I_{\mathrm{ice}}}$ and its corresponding $\sigma_{I_{\mathrm{ice}}}$ reflect the overall impact of the interaction, while $I_{pdp,ij}$ quantifies the influence of specific interactions.
@@ 5 Demonstration on analytical and engineering problems
Metrics for analytical problems were evaluated on the original function, while surrogate models were used for engineering cases where the true function was unknown. For all problems, we conducted a comparison with other sensitivity metrics (e.g., Sobol’ indices and averaged SHAP).
@@ 5.1.1 Global sensitivity analysis
First, let us compute the GSA values from Sobol’ indices, averaged SHAP, PDP, and ICE. The comparison of various GSA metrics is presented in Table 3. Notably, all metrics produce the same ranking of importance, with the most significant variable being $x_{4}$ followed by $x_{1}$ and $x_{2}$, with $x_{3}$ and $x_{5}$ are considered to be less significant. Interestingly, $I_{pdp}$ and $\mu_{I_{\mathrm{ice}}}$ provide identical importance values for $x_{3}$, $x_{4}$, and $x_{5}$. This occurs because the centered ICE curves for variables without interactions are essentially the PDP itself, leading to the mean of $I_{\mathrm{ice}}$ (i.e., $\mu_{I_{\mathrm{ice}}}$) being identical to $I_{pdp}$. Conversely, the $\mu_{I_{\mathrm{ice}}}$ values for $x_{1}$ and $x_{2}$ are higher than their $I_{pdp}$ counterparts, which is due to the interaction effects considered in the calculation of $\mu_{I_{\mathrm{ice}}}$. In this regard, if the value $I_{pdp}$ equals $\mu_{I_{\mathrm{ice}}}$, this signals that the interaction is not present. Conversely, when $\mu_{I_{\mathrm{ice}}}>I_{pdp}$, this signals that the interaction term is present since this means that the ICE curves are not exactly similar to the PDP in terms of the trend and impact. Note that this approach is similar to interpreting first-order and total order Sobol’ indices. It is also evident in this problem that the values of $S_{T}$ and $S_{F}$ are identical for $x_{3}$, $x_{4}$, and $x_{5}$, while $S_{T}$ for $x_{1}$ and $x_{2}$ are higher than $S_{F}$ for $x_{1}$ and $x_{2}$.
@@
The barplot of $I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$ for the Friedman regression problem is further shown in Fig. 2(a). Here, it can be seen that $\mu_{I_{\mathrm{ice}}}$, with the errorbar showing $\mu_{I_{\mathrm{ice}}}\pm\sigma_{I_{\mathrm{ice}}}$ provides more information than $I_{pdp}$ and $\bar{Sh}$. Again, a large vertical bar indicates a stronger effect of interactions, as evidenced in $x_{1}$ and $x_{2}$, but non-existent in other variables. Further, Fig. 2(b) depicts the violin plot of the correlation between the ICE curves and PDP. For $x_{1}$ and $x_{2}$, the overall correlations are positive, but some ICE curves deviate significantly from the PDP (indicating strong interaction), particularly those with lower correlation values. The non-zero values of $\sigma_{\rho}$ also signify that the effect of the interaction changes the trend of the ICE curves. It is worth noting that a non-zero value of the correlation between the ICE curve and PDP indicates a change in nonlinearity or trend. The $\sigma_{\rho}$ value for $x_{3}$, $x_{4}$, and $x_{5}$ is exactly zero because the trend for the ICE curve is exactly similar to its PDP. Let us now analyze the interaction GSA-values as shown in Table 4. The two-way PDP feature importance, SHAP, and second-order Sobol’ indices (see Table 4). further confirm the presence of the interaction between $x_{1}$ and $x_{2}$. In this regard, Second-order Sobol’ indices quantify the interaction effects between pairs of input variables on the output variance, while interaction SHAP values measure the specific contribution of interactions between inputs to a model’s prediction. It was observed that the estimated interaction metrics are zero or nearly zero for all cross terms but $x_{1}-x_{2}$. Note that this information should be digested together with an analysis of the visualization of PDP and ICE, as shown next.
(a) Barplot of $I_{pdp}$, $\mu_{I_{\mathrm{ice}}}$, and $\bar{Sh}$.
(b) Violin plot of ICE-based correlation values.
Figure 2: Barplot of some GSA metrics and violin plot of ICE correlation values for the Friedman regression problem.
@@ 5.1.2 Visualization
The first variable of interest is $x_{1}$, with the centered PDP/ICE and SHAP plots shown in Figs. 3 and 4, respectively. The plots for $x_{2}$ are similar, as both variables appear only in the term $10\sin(\pi x_{1}x_{2})$. The curves are colored by $x_{2}$ and $x_{4}$, where $x_{1}$ interacts solely with $x_{2}$. Notably, the centered PDP/ICE and SHAP plots reveal distinct insights: the effect of $x_{1}$ varies with the value of $x_{2}$, reflecting their interaction. For instance, when $x_{2}$ is small, the impact of $x_{1}$ is minimal, resulting in an almost flat ICE curve. However, as $x_{2}$ increases, the impact of $x_{1}$ becomes more significant, with the overall trend first increasing and then decreasing (which is due to the sinusoidal term). This effect is quantified and shown by the non-zero values of $\sigma_{I_{\mathrm{ice}}}$ and $\sigma_{\rho}$ discussed earlier. Further, it can be seen that the coloring seems to be disordered when the ICE curves for $x_{1}$ are color-coded according to $x_{4}$; which indicates that there is no interaction between how the change in $x_{4}$ affects $x_{1}$.
@@
We now address the uncertainty analysis of a wind turbine fatigue problem, where the uncertainties involve five independent random variables related to wind conditions [18]. The input variables considered are wind speed ($V_{hub}$), wind direction ($\theta_{w}$), wave height ($H_{s}$), wave period ($T_{p}$), and wind-platform direction ($\theta_{p}$) with their histograms are shown in Fig. 6. Notice that the data was normalized within $1,1]^{5}$, as provided by the original paper. The side-side tower base bending moment ($M_{x,twr}$) is selected as the output of interest due to its nonlinear behavior, evaluated using FAST, NREL’s aeroelastic code for simulating the coupled dynamic response of wind turbines [22]. Subsequently, a PCE model with $p_{max}=5$ and data-driven orthogonal polynomials was built using 2000 samples, yielding $R^{2}=0.97$ on the validation data set comprising 500 samples, which is more than sufficient for knowledge discovery purposes.